OpenStudy (anonymous):

Check my work(derivatives): let F(x) = arctan(3lnx). Find f'(e) d(3lnx) * d(arctan(3lnx)) 3/x * 1/(1 +(3lnx)^2)= 3/x(1+ 9ln^2 x) f(e) = 3/e(1+ 9ln^2 e) = 3/10e

OpenStudy (anonymous):

What's f? You specified F

OpenStudy (anonymous):

sorry, that's a technical typo.

OpenStudy (anonymous):

I tried implicit differentiation\[ \begin{array}{rcl} f(x) &=& \tan^{-1}(3\ln(x)) \\ \tan(f(x)) &=& 3\ln(x) \\ \frac{d}{dx}\tan(f(x)) &=& \frac{d}{dx} 3\ln(x) \\ \sec^2(f(x))f'(x) &=& \frac{3}{x} \\ \sec^2(\tan^{-1}(3\ln(x)))f'(x) &=& \frac{3}{x} \\ f'(x) &=& \frac{3\cos^2(\tan^{-1}(3\ln(x)))}{x} \\ \end{array} \]Then plug in \(e\) \[ \begin{split} f'(e) &= \frac{3\cos^2(\tan^{-1}(3\ln(e)))}{e} \\ &= \frac{3\cos^2(\tan^{-1}(3))}{e} \\ &= \frac{3}{e\sqrt{1^2+3^2}^2} \\ &=\frac{3}{10e } \end{split} \]